Circular No. 7766

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001hg IN NGC 4162 T. Puckett, Mountain Town, GA; and A. Sehgal, Woodinville, WA,

report the discovery of an apparent supernova (mag 17.4) on an unfiltered CCD frame taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Dec. 4.44 UT. SN 2001hg is located at α = 12^{h}11^{m}53^{s}.18, δ = +24^{o}06′49′′.4 (equinox 2000.0), which is 9′′.4 east and 35′′.3 south of the center of NGC 4162. The new object was confirmed on CCD frames taken on Dec. 5.29, but it is not present on images taken by Puckett on Mar. 12, Apr. 9, and 26 (limiting mag ∼ 20.5) or on Palomar Sky Survey images taken on 1996 May 14 (limiting mag ∼ 21.0), 1994 Jan. 15 (limiting mag ∼ 21.0), and 1950 June 10 (limiting mag ∼ 20.0).

SUPERNOVA 2001hh IN MCG –02-57-22 M. Schwartz, Cottage Grove, OR, and W. D. Li, University of Califor-

nia at Berkeley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova (mag ∼ 17.5) on unfiltered images taken with the 0.5-m Tenagra III automated telescope on Dec. 4.1 and 5.1 UT. The new object is located at α = 22^{h}36^{m}30^{s}.40, δ = −12^{o}34′08′′.0 (equinox 2000.0), which is 11′′.1 east and 7′′.1 south of the nucleus of MCG –02-57-22. A KAIT image taken on Nov. 2.1 showed nothing at this position (limiting mag ∼ 19.0).

COMET C/2000 WM_{1} (LINEAR) Further selected total-visual-magnitude and coma-diameter estimates:

Nov. 14.27 UT, 7.2, 12′ (C. S. Morris, Fillmore, CA, 10×50 binoculars; 1^{o}.1 tail); 19.95, 6.3, 18′ (R. Haver, Frasso Sabino, Italy, 10×50 binoculars); 22.91, 5.8, 20′ (B. H. Granslo, Fjellhamar, Norway, 7×50 binoculars); 30.78, 5.7, 15′ (J. J. Gonzalez, Asturias, Spain, 7×50 binoculars); Dec. 3.38, 5.4, 21′ (Y. Nagai, Yamanashi, Japan, 7×35 binoculars); 3.79, 5.3, – (Gonzalez).

COMET P/2001 MD_{7} (LINEAR) Visual m_{1} estimates: Oct. 16.53 UT, 14.1 (M. Mattiazzo, Wallaroo, S.

Australia, 0.28-m reflector); Nov. 11.46, 12.8, – (D. A. J. Seargent, Cowra, N.S.W., 0.25-m refl.); 14.44, 11.9 (K. Yoshimoto, Yamaguchi, Japan, 0.25m refl.); Dec. 2.55, 12.4 (A. Pearce, Nedlands, W. Australia, 0.41-m refl.).

2001 December 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green